A bag contains 25 chips numbered 1 to 25. If a chip is drawn randomly from the bag, what is the probability that it is 7 or 15?
Answer by math_tutor2020(3835)


Answer: 2/25 = 0.08 = 8%

Reason:
There are two chips we're after (a "7" or a "15") out of 25 total.

You could think of it like this
P(getting "7") = 1/25
P(getting "15") = 1/25
P(getting "7" or "15") = P(getting "7") + P(getting "15")
P(getting "7" or "15") = 1/25 + 1/25
P(getting "7" or "15") = 2/25